OpenStudy (anonymous):

Math help, geometry 2A

OpenStudy (anonymous):

\[14-4y \ge 38\] I got \[y \ge-6\]

OpenStudy (anonymous):

yh whats up

OpenStudy (anonymous):

ok Simplifying 14 + -4y = 38 Solving 14 + -4y = 38 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '-14' to each side of the equation. 14 + -14 + -4y = 38 + -14 so after that what do you do while combining the terms

OpenStudy (anonymous):

well ill just show you Combine like terms: 14 + -14 = 0 0 + -4y = 38 + -14 -4y = 38 + -14 Combine like terms: 38 + -14 = 24 -4y = 24 Divide each side by '-4'. y = -6 Simplifying y = -6

OpenStudy (anonymous):

so you got it right @kelliandco

OpenStudy (anonymous):

I didn't.

OpenStudy (anonymous):

what happened